OpenStudy (anonymous):

36a^2/4a^2-8a+4 * 4a-4/6a= I got 8a+25/3 is this correct

jhonyy9 (jhonyy9):

9-8a+16a-2/3a = 8a +9 -2/3a = (24a +27 -2a)/3 =(22a +27)/3